2015-Calendar-By-Week-4.jpg

Week14:

  • Mar30:
DQM validation : 740_pre9 vs 740pre8_root6 - done Wg-kafactor: code preparation for gen-level analysis eTnP:

Week13

  • Mar25:

LHC restart today not possible: http://press.web.cern.ch/

Week11

  • Mar14-15:
GTC request for HcalElectronicsMap_v7.05_hlt tag,

  • Mar19:
Wg: use SCRiso now, fit phoSCRChIso distribution without phoPFChIso cut, apply efficiency on phoSCRChIso rather than phoPFChIso which can be taken from signal template, Ming is fine and suggest "it would be just sufficient to check the SF difference between two sets of ID (phoPFChIso+other variables vs phoSCRChIso+other variables) with Z->mumugamma events by simple counting. You can apply the difference to correct the recommended POG SF and add a systematic uncertainty." some comparisons between these two isolations in these slides: http://lovedeep.web.cern.ch/lovedeep/WORK15/SCRVsNominal_Iso_Mar19.pdf (signal is from Wg and background is Wjets simulated sample). From these slides, "SCR-ChIso < 2.0 in EB, EE) gives optimized values, but if we want same signal efficiency, then we can keep 1.5 in EB, and 1.2 in EE. YM: prefer to have the same efficiency.

Week11

  • Mar9:
DQM-offline, validated 740_pre8 Vs 740_pre6 (hist name changed) and 740_pre8_root6 vs 740_pre8

News: latest menu with rates: http://tinyurl.com/qhngww4 (https://hypernews.cern.ch/HyperNews/CMS/get/susy/2003.html)

HCAL problems:

RooFit for peak fit: https://twiki.cern.ch/twiki/pub/CMS/SWGuideCMSDataAnalysisSchoolStatistics2015Bari/exercise_0.C

Week10

  • Mar2:
Maxim Azarkin - new person for validation of HCAL conditions, AlCaDB meeting (8-10AM): my 1st: https://indico.cern.ch/event/368329/ couldn't learn anything, 11AM- group meeting - get AN for zgg etc., ST templates looks bad, sidebands give huge variation in sig yields. and about egamma queries plan to get PAS for EXO-14-015 - need to remake abcd plots with cms prel.

  • Mar3:
ST new AN for zgg, TOTW poster, Wg - iso templates

  • Mar4:

Clemencia Mora wrote to me: I updated the offline and hlt tags and created the MC tag for the updated QIE data. I also added a small description to the twiki: https://twiki.cern.ch/twiki/bin/viewauth/CMS/HcalQIEDataTags2011 Wg meeting - plan AN, slides here: https://indico.cern.ch/event/378500/contribution/2/material/slides/0.pdf

  • Mar5:

HCAL has something bad,

Clemencia appended a new IOV to the existing tag (in use) - problem she describes "I naively trusted in that the experts that sent the new table to me knew what was/wasn't there in the input" -

Offline Run Manager ORM this week - roberto castello not happy

salvt mention "Andre Holzner frm P.5 and quickly figured out this tag has got broken IOV last night (~2am CET): there were 22 lines (ZDC readouts) "cut off"... HLT guys have been notified, as well as HCAL management. Clemencia has fixed it in new IOV 236955.

HLT streams are crashing without a fix (appended 22 ZDC lines)...

Marco Musich report an error related to an upload to the "HcalQIEData_NormalMode_v7.00_hlt" tag done today at 1.51 PM from Clemencia

TODO: DQM validation - deadline is March 9 and 10 for 740_pre8 & 740_pre8root6

AN-15/55 for Wg

  • Mar6:

HCAL DPG: https://indico.cern.ch/event/379195/ (7AM-9AM), svn co for AN-15-54/55,

Week9

  • Feb23:
YM "L4 convener in the HCAL calibration - to be responsible for the conditions and be a liaison to the Operations" group meeting - 11-12PM DQM-offline service task - step back by May-end - will get 1 month credit meet jordan to learn Hcal l4 responsibilities - 3PM - read twikis etc. sent by email

  • Feb24 - Tue - Andrew's new student would work on DQM - from mid-june, Peggy about M.Leave, YM forwarded QIE update email from Clemancia

  • Feb25:
Wg meeting : https://indico.cern.ch/event/376944/ - I hav summary of Sieie templates, create Wg repo in my github, go back to SCR iso templates - this time no cut on ChHadIso,

HCAL task learning:

IOV=233723 since that has appropriate ZDC gains?

  • Feb26: Thu: Salavt says "for hlt/_express flavors last IOV is automatically assigned by the uploading system from the Run registry to set it to the _most recent taken run. That's why it comes "233722, 235714" and not 233722, 233733"

my task is to submit tags with account info. to GTc tool, what about queues - ask ppd-conv-alcadb., subscribed hn-cms-alca and "cms-project-alcacontacts"

  • Feb27:
Hcal calibration meeting (7-9am), GTC and Alcadb email etc. - wait now for salavt, start again with Wg - move back to SCR templates (reading AN-12-485).

Marco M says HcalGains_v2.08_hlt tag already there in last GT in the queue (GR_H_V54). Salavt: it has got an addional IOV. Marco: if the tag is there, and you just appended an IOV, you don't need to re-queue. It should be announced on the alca hypernews (hn-cms-alca@cern). Salavatore di Guida: in principle, any update of the conditions in a data processing environment should undergo validation. I'll do this? Salaavt: I don't think we're in position to present the entire summary of changes...important summary in email

Week8

  • emu-for S.T., Wg for me (weekend)
  • Feb16: fixed AN-14-148, EXO-14-015 text, TOTW poster, group meeting, l*-twiki, Wg-AN
  • Feb17 - Tue*: Wg-AN+slides - template method -jets faking as photons - SCR-CHIso (SB variation etc.), pass TnP tools to Giovanni
  • Feb18 - Wed*: https://indico.cern.ch/event/375292/contribution/2/material/slides/0.pdf (Wg meeting), try to check with Sieie templates, if I get similar results. Ilya suggested don't say fake rate.
  • Feb19 - Thurs - katya preparing Wg quick update slides for smp-vv, I on Wg-Sieie templates
  • Feb20 - Fri - smp-vv - 9-11AM - https://indico.cern.ch/event/370169/

Week7

  • split lhe file with 10K events into 1K files, so I have 10 events in each lhe file. Feed these 1K lhe files to 1K cfg files to run GENtoRECO step and submit crab. link to my git area (weekend)
  • Feb9:
TOTW poster, group meeting, l*analysis: prepare slides for reply to hn (lstar to CI decays). Approval comments answered here: http://lovedeep.web.cern.ch/lovedeep/WORK15/2l2j_CIanalysis_inreply.pdf , DQM: https://hypernews.cern.ch/HyperNews/CMS/get/relval/3395/51.html - 740_pre6Vs740_pre5, https://hypernews.cern.ch/HyperNews/CMS/get/relval/3363/72.html - 740_pre5Vs740_pre2 (MON)
  • Feb10:
Got comments on l* slides from E.P and M.M only. Steve showed interest in this work...sent to him too. Sent hn https://hypernews.cern.ch/HyperNews/CMS/get/EXO-14-015/25/2/1.html to EXO-Conv. Steve told there was discussion on this topic in EXO-cordination meeting. TOTW. (TUE)
  • Feb11:
Wg meeting, Zgg, CoffeeHour, No reply to l* hn, YM-no need to apply for hcall4-positn, Wg-AN (WED) l* approved! https://hypernews.cern.ch/HyperNews/CMS/get/EXO-14-015/27/1.html, I will add an appendix in AN-14-148 and 2-3 sentences in paper for MC-based cross-checks. (FRI)

-- LovedeepKaurSaini - 2015-02-11

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2015-03-30 - LovedeepKaurSaini
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback